PLEASE CHECK MY ANSWER.
Find the equation of the inverse of f(x)=x^2+2. If the inverse is a function, write the result using the notation f^-1(x).
I got answer of y=x^2+2.. Is this right?

Answer by ewatrrr(24785)

y=x^2+2 to find Inverse put x in place of y
x = y^2 + 2
± sqrt(x-2) = y
Inverse is not a function (y-value with 2 x-values)